OpenStudy (amy0799):

If f(x)=x^2-x and g(x)=1/x, and h(x) is defined with the expressions below, arrange the values of h'(1) from smallest to largest. 1. h(x) = f(x)-g(x) 2. h(x) = f(x)g(x) 3. h(x) = f(x)/g(x)

OpenStudy (amy0799):

do you know how to do this? @IrishBoy123

OpenStudy (dinamix):

h'(1) = f'(1)-g'(1) u must find f'(x) and g'(x)

OpenStudy (dinamix):

@amy0799

OpenStudy (irishboy123):

where are you stuck?

OpenStudy (dinamix):

for 2- h'(1) =f'(1)g(1)+g'(1)f(1) 3- h'(1) =( f'(1) g(1) - f(1)g'(1) )/ (g(1))^2 this the answer @IrishBoy123 right lol

OpenStudy (amy0799):

h(x) = f(x)-(x) =\[x^{2}-x-\frac{ 1 }{ x }\] \[=2x-1-\frac{ 1 }{ x^{2} }\] Is this correct?

OpenStudy (irishboy123):

not quite last term?

OpenStudy (amy0799):

the last term is wrong?

OpenStudy (irishboy123):

\[-\frac{ 1 }{ x } = - x^{-1}\]

OpenStudy (irishboy123):

differentiate that again

OpenStudy (amy0799):

\[2x-1-\frac{ 1 }{x^{-2} }\]

OpenStudy (dinamix):

2x-1-(-1/x^2) = f'(x)-g'(x)

OpenStudy (dinamix):

@amy0799 its like my answer see it

OpenStudy (irishboy123):

\[\frac{d}{dx} (\frac{1}{x}) = \frac{d}{dx}(x^{-1}) = -x^{-2} = -\frac{1}{x^2}\] the sign changes :p

OpenStudy (amy0799):

y wouldn't it be x^-2?

OpenStudy (irishboy123):

\[2x-1-(-1/x^2) = 2x-1+1/x^2\]

OpenStudy (dinamix):

x^-2 = 1/x^2 u are right but we dont need it @amy0799

OpenStudy (irishboy123):

you had it right with our first go except you got the sign wrong on last term

OpenStudy (dinamix):

me ? @IrishBoy123 u spoke with me

OpenStudy (dinamix):

lol

OpenStudy (irishboy123):

no @dinamix :p i was addressing @amy0799

OpenStudy (amy0799):

h(x)=f(x)g(x) \[=x^{2}-x(-\frac{ 1 }{ x^{2} })+\frac{ 1 }{ x }(2x-1)\] is this right so far?

OpenStudy (dinamix):

@amy0799 yup this

OpenStudy (amy0799):

\[\frac{ -x-1 }{ x }+\frac{ 2x-1 }{ x }\]

OpenStudy (amy0799):

would this simplify to \[\frac{ -x-2 }{ x }?\]

OpenStudy (dinamix):

-x+1/x @amy0799

OpenStudy (dinamix):

not -x-1/x

OpenStudy (irishboy123):

i'm afraid i am not following what you have done so i shall leave it to you and @dinamix

OpenStudy (dinamix):

ok @IrishBoy123 i help him ty anyway for coming

OpenStudy (amy0799):

-1+(-1)=-2 I dont understand how you got -1

OpenStudy (dinamix):

\[(x^2-x)(-1/x^2) = -1+\frac{ 1 }{ x } = \frac{ -x+1 }{ x}\]

OpenStudy (dinamix):

u made mistake @amy0799 u find it ?

OpenStudy (dinamix):

f(x)g(x) = 1 must find it like @amy0799 so check and tell me

OpenStudy (amy0799):

\[(x^{2}-x)-\frac{ 1 }{ x ^{2} }=-\frac{ x ^{2} -x+1}{ x ^{2} }\]

OpenStudy (amy0799):

is this how you set it up?

OpenStudy (dinamix):

\[\frac{ -x^2+x }{ x^2 } \]

OpenStudy (dinamix):

its like that @amy0799

OpenStudy (amy0799):

oh ok

OpenStudy (dinamix):

so what u find f(x)g(x) = i want your answer

OpenStudy (amy0799):

h(x)=f(x)/g(x) \[=\frac{ \frac{ 1 }{ x(2x-1)-(x ^{2}-x)-\frac{ 1 }{ x ^{2} } } }{ (\frac{ 1 }{ x })^{2} }\]

OpenStudy (amy0799):

is this setup right?

OpenStudy (amy0799):

i messed up, it's 1/x

OpenStudy (dinamix):

u make some mistake

OpenStudy (dinamix):

\[\frac{ \ ( \frac{ 1 }{ x})( 2x-1) -[\frac{ -1 }{ x^2} (x^2-x)]} { \frac{ -1 }{ x^2 } }\]

OpenStudy (dinamix):

@amy0799

OpenStudy (dinamix):

its not - 1 only 1

OpenStudy (dinamix):

in demonst

OpenStudy (irishboy123):

where did this come from?!?!?! https://gyazo.com/b501b2ad7a4061eab75c1cf7645b4f67 \[h(x) = \frac{f(x)}{g(x)} = \frac{x^2 - x}{\frac{1}{x}} = x(x^2 - x)\] now, expand and differentiate. that's all there is to it and did we ever come up with an answer to pt 1?

OpenStudy (dinamix):

\[\frac{ (2x-1+x-1)x^2 }{ x }\]

OpenStudy (dinamix):

we calcul h'(x)

OpenStudy (dinamix):

by using f'(x) and g'(x) thats only @IrishBoy123

OpenStudy (irishboy123):

and what exactly is this all about?!?!?!?! https://gyazo.com/d2c0f8f760ddd3e17b317893a135989d \[h(x) = f(x)g(x) = (x^2 - x). \left(\frac{1}{x}\right) = x - 1\] \[h'(x) = 1\] done.

OpenStudy (irishboy123):

if you want to collate your solutions, i am happy to come back later and check them @freckles

OpenStudy (dinamix):

we have 2 method @amy0799

OpenStudy (amy0799):

\[\frac{ (3x-2)x ^{2} }{ x }\] can this simplify further?

OpenStudy (dinamix):

yup

OpenStudy (dinamix):

h'(x) for the last one

OpenStudy (dinamix):

h'(x) for function f(x)/g(x) = (3x^2-2x)

OpenStudy (dinamix):

i mean h'(x) =(3x^2-2x) not f(x)/g(x) did u understand

OpenStudy (dinamix):

all this time we calcul h'(x) not h(x) @amy0799

OpenStudy (dinamix):

did u understand mate what we are do

OpenStudy (amy0799):

woudn't it be \[\frac{ (3x ^{3}-2x ^{2}) }{ x }?\]

OpenStudy (dinamix):

x^2/x = x right @amy0799 and its same

OpenStudy (dinamix):

and u have devide by x

OpenStudy (amy0799):

oh ok. so 3x^2-2x is the answer?

OpenStudy (dinamix):

yup

OpenStudy (amy0799):

the question also asked arrange the values of h'(1) from smallest to largest. do i plug in 1 to x?

OpenStudy (dinamix):

and this h'(x) for function f(x)/g(x) , u understand what was do

OpenStudy (dinamix):

we find all h'(x) for this fuction f(x)-g(x) and f(x)g(x) and f(x) /g(x)

OpenStudy (dinamix):

right ?

OpenStudy (amy0799):

right, so what does arrange the values of h'(1) from smallest to largest mean?

OpenStudy (dinamix):

h'(1) = f'(1)-g'(1) = 2(1)-1+1(1^2) = ?

OpenStudy (dinamix):

thats only

OpenStudy (amy0799):

=2 right?

OpenStudy (dinamix):

1- h'(1) = 2 yup

OpenStudy (amy0799):

for f(x)/g(x) 3(1)^2-2(1)=1

OpenStudy (dinamix):

yup this

OpenStudy (dinamix):

and for fuction f(x)g(x) ; h'(1) = ?

OpenStudy (amy0799):

so from smallest to largest it would be f(x)g(x), f(x)/g(x), f(x)-g(x) f(x)g(x)=1

OpenStudy (dinamix):

not f(x)g(x) =1 its f'(x)g'(x) = 1

OpenStudy (dinamix):

@amy0799

OpenStudy (dinamix):

why u make this mistake ?@amy0799

OpenStudy (amy0799):

what mistake?

OpenStudy (dinamix):

look up what i write u

OpenStudy (dinamix):

not f(x)g(x)= 1 its f'(x)g'(x) = 1

OpenStudy (amy0799):

oh i forgot to put in the '

OpenStudy (dinamix):

are u sure u understand now .@amy0799

OpenStudy (amy0799):

yes so the order goes f'(x)g'(x)=1 f'(x)/g'(x)=1 f'(x)-g'(x)=2 ?

OpenStudy (dinamix):

f'(1)g'(1)=1 f'(1)/g'(1) =1 f'(1)-g'(1)=2 not like u said

OpenStudy (dinamix):

@amy0799

OpenStudy (amy0799):

oh ok thank you. can you help me with another problem?

OpenStudy (dinamix):

ok but fast

OpenStudy (amy0799):

if f(x)=x^3+2x^2-5x-1 find \[\lim_{h \rightarrow 0}\frac{ f'(-3+h)-f'(-3) }{ h }\]

OpenStudy (dinamix):

\[\lim_{h \rightarrow 0}\frac{ f(-3+h)-f(-3) }{ h}\]

OpenStudy (dinamix):

@amy0799 not like u said its not f'(-3+h) -f'(-3) look good how i write it